OpenStudy (anonymous):

MEDAL!!! 3/4 divided by 5/18 = ? Find the quotient and simplify.

OpenStudy (astrophysics):

\[\huge \frac{ \frac{ 3 }{ 4 } }{ \frac{ 5 }{ 18 } } \implies \frac{ 3 }{ 4 } \times \frac{ 18 }{ 5 }\] when dividing two fractions you take the second fraction, flip it and multiply.

OpenStudy (anonymous):

2.7?

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

Hey sorry, I was away...but yeah that's right.